OpenStudy (johan14th):

How to find the derivative of (x)(e^x) -(e^x) + (2)

OpenStudy (anonymous):

Let y=x e^x-e^x+2 \[\frac{ dy }{ dx }=(x e^x+1*e^x)-e^x=xe^x\]

OpenStudy (johan14th):

Could please you simplify the process you did?